If 149.1 g of KCI is dissolved in water to make 500 ml solution, what is the molarity of the solution?

1
Expert's answer
2022-04-08T03:00:04-0400

Solution:

The molar mass of KCl is 74.55 g/mol

Therefore,

Moles of KCl = (149.1 g KCI) × (1 mol KCl / 74.55 g KCl) = 2.0 mol KCl


Liters of solution = (500 mL) × (1 L / 1000 mL) = 0.5 L


Molarity of KCl solution = Moles of KCl / Liters of solution

Molarity of KCl solution = (2.0 mol) / (0.5 L) = 4.0 M

Molarity of KCl solution = 4.0 M


Answer: The molarity of KCl solution is 4.0 M